OpenStudy (anonymous):

Solve for x: |x + 2| + 16 = 14 (2 points) x = -32 and x = -4 x = -4 and x = 0 x = 0 and x = 28 No solution

OpenStudy (xapproachesinfinity):

no solution

OpenStudy (anonymous):

|x+2|+16-16=14-16 =>|x+2|=-2 removing magnitued give rise to +- so =>x+2=+-2 =>either x+2=-2 or x+2=2 now solve both of these for x..

OpenStudy (anonymous):

is the answer really no solution

OpenStudy (xapproachesinfinity):

abolute value cannot produce negative numbers therefore no solution

OpenStudy (anonymous):

true true